Suppose that a certain college class contains 54 students. Of these, 27 are sophomores, 33 are economics majors, and 10 are neither. A student is selected at random from the class.
(a) What is the probability that the student is both a junior and a psychology major?
(b) Given that the student selected is a junior, what is the probability that he is also a psychology major?
Write your responses as fractions.

Answer by stanbon(75887)

Suppose that a certain college class contains 54 students. Of these, 27 are juniors, 33 are psychology majors, and 10 are neither. A student is selected at random from the class.
----
Draw a Venn Diagram with two intersecting circles: J and E
---
Let the intersection contain x students.
Equations:
x + J = 27
x + P = 33
J + x + P + 10 = 54
---
J = 27-x
P = 33-x
-----
Substitute and solve for "x":
27-x + x + 33-x = 44
-x = -16
x = 16 (# who are juniors and psych majors)
=====================================================
(a) What is the probability that the student is both a junior and a psychology major? 16/54
--------------------
(b) Given that the student selected is a junior, what is the probability that he is also a psychology major?::: 21/33
--------------------------------------------
Write your responses as fractions.
===========
Cheers,
stan H.